A cord pulling a cart up along a frictionless plan
Answer #1

Since there is no friction, only two forces are acting along incline, one is tension which is pulling up the plane and component of gravitational force, pulling down the plane.

The component of gravitational force along incline is mg*sin(15)

Net force = T- mg*sin15o

Net force = 25 - 5*9.81*sin(15o)

Net force = 12.305 N up the plane

1) Acceleration = force/mass = 12.305/5

a = 2.46 m/s

2) Normal force is component of weight

N = mg*cos(15)

N = 47.38 Newtons